OpenStudy (anonymous):

Mars is about 142 million miles from the sun. The earth is about 93,000,000 miles from the sun. How much farther from the sun is Mars than the earth? awnser in scientific notation

OpenStudy (anonymous):

ALRIGHT brah. So I'm sure you know how to find out how much farther away they are. That's some easy stuff. Basically take the bigger number and subtract it by the smaller number there's your answer. I assume the part that's throwing you off is the scientific notation.

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

To do that pellet. Like say we have the the number 3,300,000 right? Well in scientific notation all numbers are written with two places like this. 2.2 1.1 4.4 etc So with the number 3,300,000 it would be 3.3 now after you do all that stuff you have to multiply it by 10 so 3.3 * 10 Now after you do that you gotta find raise 10 to a power to make it equal to a million. 10^1 = 10 10^2 = 100 10^3 =1000 10^4 = 10,000 10^5 = 100,000 10^6 = 1,000,000 note that for each time you raise ten to a power you're adding a zero each time. Now to do this for 3,300,000 You see 3.3 times what number = 3.3 million? You see that 10^6 = 1,000,000 so 3.3 times a million is 3.3 million so 3.3 million in scientific notation is 3.3 * 10 ^6

OpenStudy (anonymous):

ok thanks bro

OpenStudy (anonymous):

Yeah. So 400 in scientific notation would 4.0 * 10^2 4000 would be 4.0 * 10^3 etc etc. You're trying to find scientific notation of 49 million so it would be 4.9 * 10 ^ 7